OpenStudy (anonymous):

if $300 is deposited in an account and it increases to $450 in eight years compounded quarterly. find the rate of interest to the nearest tenth

OpenStudy (amistre64):

450 = 300(1+ r/4)^(8r)

OpenStudy (anonymous):

and how do i solve that @amistre64

OpenStudy (amistre64):

with cunning and style

OpenStudy (amistre64):

46.28%

OpenStudy (amistre64):

divide 300 take logs; and use rules of logs to manipulate it

OpenStudy (amistre64):

or graph it and find the zero

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can you show me step by step

OpenStudy (amistre64):

i can try lol

OpenStudy (anonymous):

ok thanks lol

OpenStudy (anonymous):

@ amistr64 can i ask you a question when you done helping them

OpenStudy (amistre64):

its easier on paper... 450/300= 3/2 3/2 = (1 + r/4)^(8r) log(3/2) = log(1+ r/4)^(8r) log(n) = 8r log(1+ r/4) log(n) ---------- = 8r log(1 + r/4)

OpenStudy (anonymous):

so problems like this have to be solved with a calculator?

OpenStudy (amistre64):

log(3) - log(2) -------------- = 8r log(4+r) - log(4) A ---------- = 8r log(4+r) - B log(4+r) - B = A/8r etc...

OpenStudy (amistre64):

it canbe done by hand, but its alot of window dressing to get it to the right spot

OpenStudy (anonymous):

and the way you doing it here is by hand or calculator

OpenStudy (anonymous):

@amistre are you still there

OpenStudy (amistre64):

by hand; which is why i left it alone really; :)

OpenStudy (anonymous):

oh ok